# Learning the bulk and interfacial physics of liquid-liquid phase separation with neural density functionals

### v2

@@ Abstract
where $T^{\ast}=k_{B}T/\epsilon$ is scaled temperature, $b$ is a constant, the critical concentration is $1/2$ per symmetry, and the exponent $\beta$ should neither be confused with inverse temperature nor with the liquid phase in which species 1 is the majority component. Because of the Ising-like symmetry there is no need to include a linear contribution in Eq. (3). The three-dimensional Ising value is $\beta=0.32630(22)$ Ferrenberg, Xu, and Landau (2018), but here the exponent $\beta$ is treated as a free fit parameter. Following Ref. Sammüller, Schmidt, and Evans, 2025 we exclude data very close to the critical $\lambda$ point, using only coexisting densities for $k_{B}T/\varepsilon\leq 1.06$, and find from the fitting $T_{\lambda}^{\ast}=1.118$ and $\beta=0.329$ for this particular total density, see Fig. 3. $T_{\lambda}$ obtained from the fitting procedure is about $3.5\%$ lower than the value obtained from direct evaluation of the neural functional by tracing the $\beta\gamma$ binodal, i.e. calculating the merging of coexisting density profiles or, equivalently, from the vanishing of the second-derivative of the grand potential.
Our framework allows us to access interfacial tensions via expressing the excess free energy difference $\Delta F_{\rm exc}$ as a functional integral,
@@
We return now to the full phase diagram obtained from considering density profiles such as those in Fig. 2. Analysis of the plateau values of all stable fluid interfaces allows us to construct the entire fluid phase diagram, shown in the top panel of Fig. 4, and plotted as a function of the total density $\rho_{1}+\rho_{2}$ and temperature $T$. We have checked that for $T>T_{\rm CEP}$ the results obtained for coexisting densities are numerically consistent with those obtained via the Maxwell construction based on bulk functional integration Sammüller et al. (2023). The representation in Fig. 4 makes apparent the density jump at gas-liquid coexistence. Recall that above the CEP the liquid is in a mixed state; both species have equal concentration. Below the CEP the liquid is demixed, as discussed above, and the concentration jump (see Fig. 3) is ‘collapsed’ in the representation of Fig. 4. As described above, the $\lambda$ line is calculated from the locus of vanishing second derivative of the grand potential $\Omega$ with respect to composition $X_{1}$, readily facilitated via automatic differentiation of the neural functional Sammüller et al. (2023). Whilst our results for the gas-liquid binodal agree closely with the independent simulation data of Wilding Wilding (1997), our $\lambda$ line displays a small offset, typically about $7\%$ in density, with respect to the corresponding simulation results.
@@
In conclusion, we have investigated the bulk fluid phase behaviour and associated interfacial phenomena in a symmetrical LJ mixture with a specific choice of weakened cross attraction. Whilst imposing symmmetry serves conveniently to reduce the parameter space, the model already incorporates a wealth of physics that is challenging to describe within a machine learning context: important is the lack of a priori symmetry breaking. We adapted previously developed local learning Sammüller et al. (2023) by thermally training Sammüller, Schmidt, and Evans (2025) a single neural network to represent the pair of species-resolved one-body direct correlation functionals simultaneously. This strategy is consistent with the formal generation of $c_{1}^{(i)}({\mathbf{r}};[\rho_{1},\rho_{2}],T)$ via functional differentiation from a single, underlying, excess free energy functional.
@@ Acknowledgements.
The data that support the findings of this study are openly available  [^1] Code and models available at https://github.com/SilasRobitschko/SWNeural, complete simulation dataset available at https://doi.org/10.5281/zenodo.15777201.

### v3

@@ Abstract
[…]
Our framework allows us to access interfacial tensions via expressing the excess free energy difference $\Delta F_{\rm exc}$ as a functional line integral,
@@
We return now to the full phase diagram obtained from considering density profiles such as those in Fig. 2. Analysis of the plateau values of all stable fluid interfaces allows us to construct the entire fluid phase diagram, shown in the top panel of Fig. 4, and plotted as a function of the total density $\rho_{1}+\rho_{2}$ and temperature $T$. We have checked that for $T>T_{\rm CEP}$ the results obtained for coexisting densities are numerically consistent with those obtained via the Maxwell construction based on bulk functional integration Sammüller et al. (2023). The representation in Fig. 4 makes apparent the density jump at gas-liquid coexistence. Recall that above the CEP the liquid is in a mixed state; both species have equal concentration. Below the CEP the liquid is demixed, as discussed above, and the concentration jump (see Fig. 3) is ‘collapsed’ in the representation of Fig. 4. As described above, the $\lambda$ line is calculated from the locus of vanishing second derivative of the grand potential $\Omega$ with respect to composition $X_{1}$, readily facilitated via automatic differentiation of the neural functional Sammüller et al. (2023). Whilst our results for the gas-liquid binodal agree closely with the independent simulation data of Wilding Wilding (1997), our $\lambda$ line displays a small offset, typically about $10\%$ in density, with respect to the corresponding simulation results.
@@
In conclusion, we have investigated the bulk fluid phase behaviour and associated interfacial phenomena in a symmetrical LJ mixture with a specific choice of weakened cross attraction. Whilst imposing symmmetry serves conveniently to reduce the parameter space, the model already incorporates a wealth of physics that is challenging to describe within a machine learning context: important is the lack of a priori symmetry breaking. A single neural network was used to represent the pair of species-resolved one-body direct correlation functionals simultaneously. This strategy is consistent with the formal generation of $c_{1}^{(i)}({\mathbf{r}};[\rho_{1},\rho_{2}],T)$ via functional differentiation from a single, underlying, excess free energy functional.
Training the neural functional is based on adapting both the local learning Sammüller et al. (2023) and thermal training Sammüller, Schmidt, and Evans (2025) schemes that were previously developed for single-component systems. Using these methods directly for the present binary mixture allows one to learn solely on the basis of the available simulation data. We have found that small artifacts can arise thereby in the predictions of the precise shape and location of the gas-liquid binodal when using the above schemes without modification. These artifacts are prevented by including $L_{2}$ regularization [^1] Using $L_{2}$ regularization during training is a standard practice in machine learning which penalizes large values in the neural network parameters. This supports the generation of less complex (‘smooth’) models, which is particularly suitable for the functional mapping $c_{1}^{(i)}(\mathbf{r};[\rho_{1},\rho_{2}],T)$. in the training protocol, which we have hence adopted for the neural functional that was used to generate all numerical results shown in this work. The two different neural models trained with and without $L_{2}$ regularization, together with a comparison of the corresponding numerical predictions, are available in Ref. Zen, . In future work, it would be interesting to investigate physics-informed regularization, e.g. by incorporating exact statistical mechanical sum rulesHermann and Schmidt (2021) to provide additional constraints that are to be optimized during training.
@@ Acknowledgements.
The data that support the findings of this study are openly available Zen .